how do i solve y=-x
also how do i solve x=3y?

Found 2 solutions by Theo, Gogonati:
Answer by Theo(13342)

your equation is y = -x

your dependent variable is y and your independent value is -x.

you pick values of x and then you find values of y that correspond to that.

when x = 1, y = -1
when x = 2, y = -2
when x = 3, y = -3
when x = 0, y = 0
when x = -1, y = 1
when x = -2, y = 2
etc.

your equation is now x = 3y

your dependent variable is now x and your independent variable is now y

you would select values of y and then you would solve for values of x.

when y = 1, x = 3*1 = 3
when y = 2, x = 3*2 = 6
when y = -5, x = 3*(-5) = -15
etc.

with the equation x = 3y, if you want to make x the independent variable and you want to make y the dependent variable, you would need to switch the equation around.

you would do that as follows:

x = 3*y

divide both sides of the equation by 3 to get:

x/3 = y

flip the sides of the equation to get:

y = x/3

you have now made x the independent variable and y the dependent variable.

when x = 3, y = 3/3 = 1
when x = 6, y = 6/3 = 2
when x = -15, y = -15/3 = -5



Answer by Gogonati(855)

Solution: For y=-x all pairs of opposite numbers are the solution. (-2,2), -5,5),
For x=3y the solution are pairs (n,3n) where n is a real number i.e
(1,3), (-1,-3),(2,6), (-2,-6) e.c.t.